OpenStudy (anonymous):

Find the value of I = integral of ((4arcsin(x))/sqrt(1-x^2))

OpenStudy (anonymous):

\[I = \int\limits_{0}^{1/2} \frac{ 4\sin^{-1} x }{ \sqrt{1-x^2} }\]

OpenStudy (anonymous):

Try a substitution. If I recall correctly, the derivative of arcsine gives you the expression in the denominator.

OpenStudy (anonymous):

It worked thanks a lot!